2x-3y=-12 what is the x slope and the y intercept
Answer by ewatrrr(24785)


Hi,
Note: the standard slope-intercept form for an equation of a line is y = mx + b
where m is the slope and b the y-intercept.
2x-3y=-12 |Solving for y to put into the slope-intercept form
y = (2/3)x + 4
slope m = 2/3 and the y-intercept is Pt(0,4)
Pt(-3,2) and pt(3,6) also on the line according to the slope